OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, can be written as 7sqrt(a).

OpenStudy (anonymous):

@jim_thompson5910 , can you help again ?

jimthompson5910 (jim_thompson5910):

Think of any number you want

jimthompson5910 (jim_thompson5910):

Which number comes to mind

OpenStudy (anonymous):

10 ?

jimthompson5910 (jim_thompson5910):

How about we double that to 20 (since that works better)

jimthompson5910 (jim_thompson5910):

now let's use the variable x

jimthompson5910 (jim_thompson5910):

So let's simplify \[\Large \sqrt{20x}\] ----------------------- \[\Large \sqrt{20x}\] \[\Large \sqrt{4*5x}\] \[\Large \sqrt{4}*\sqrt{5x}\] \[\Large 2\sqrt{5x}\] ----------------------- So \[\Large \sqrt{20x}\] simplifies to \[\Large 2\sqrt{5x}\]

jimthompson5910 (jim_thompson5910):

Do you see how I'm doing this?

OpenStudy (anonymous):

Yes, thanks

jimthompson5910 (jim_thompson5910):

alright great

jimthompson5910 (jim_thompson5910):

yw